Please help:
Find the measure of each angle: Isosceles Triangle:
Angle A: 4x + 2y
Angle B: 9y - 2
Angle C: 10x + 10
Angles B and C have the same measurement.

Answer by swincher4391(1107)

By our Triangle Angle-Sum Theorem, 4x+2y + 9y -2 + 10x+10 = 180
We also know that 9y -2 = 10x + 10, so solve for y.
9y = 10x + 12
y = (10x+12)/9
4x + 2(10x+12)/9 + 9(10x+12)/9 - 2 + 10x + 10 = 180
4x + (20/9)x + 24/9 + 10x + 12 -2 + 10x + 10 = 180
(236/9)x + 68/3 = 180
Multiply everything by 9 to get:
236x + 204 = 1620
236x = 1416
x = 6
10x + 10 = 9y-2
10(6) + 10 = 9y -2
72 = 9y
y = 8
So, angle A: 4(6) + 2(8) = 40
angle B: 9(8) -2 = 70
angle C: 10(6) + 10 = 70
40,70,70